# Chrono[Uncaptioned image]: A Simple Blueprint for Representing Time in MLLMs

### v5

@@ 4.1.2 Design of timestamps
A central question for this work is how MLLMs can best represent and reason about timestamps, as they have to be interpreted in the input and predicted in the output. We compare different options for representing the timestamps and their impact on the final downstream model performance. We explore using relative positions w.r.t the video duration vs. representing each timestamp as absolute time. We compare representing these as decimal numbers (absolute “79.9” seconds, relative “0.40”) vs. integers (absolute “80” seconds, relative “40”). In these experiments, the format of the timestamps in model input and output is always consistent. In the case of relative positions, we post-process the output to yield an absolute value to compare to the ground truth for the final evaluation.
@@
Inspecting the Chrono-BLIP ablations in Table 1 (b), we observe that representing the timestamps in decimal form results in lower performance for both absolute and relative form, see rows (1,2) vs. rows (3, 4). We hypothesize this is the case because of how decimal numbers are tokenized. Although using decimal timestamps allows for higher temporal resolution, we hypothesize that this lower performance can be attributed to limitations related to the tokenization of decimal numbers. Depending on the tokenizer, a decimal number such as “79.9” or “42.05” can be divided into a different number of tokens each, whereas integers (up to a certain number) are tokenized as a single token. Nevertheless, it is important to note that this behavior varies across tokenizers and can, therefore, be subject to variance. Further, we observe that absolute position in seconds yields better performance than the relative representation when using integer precision. This shows the importance of designing the input and outputs with the representation of the MLLM in mind.
@@
When analyzing the trends for Chrono-GPT, we find that these trends are not as evident. We observe that there are a lot of variances when evaluating Chrono-GPT on the downstream tasks and that it seems, when inspecting only the first four rows, that GPT-4o prefers decimal precision. We hypothesize that the GPT-4o tokenizer is more advanced and suitable for processing numbers, having a more consistent tokenization scheme. Yet, when comparing integer and decimal precision in the interleaved setting, we find that GPT-4o prefers integer precision (see Appendix Section C.2).
@@
Finally, we explore the design choice of ordering frame and time tokens. In rows (1) through (4), we concatenate all frames together, followed by a concatenation of the timestamps, which are each separated by a separator token (“$>$”). In rows (5) and (6), we observe the benefit of interleaving frame tokens with time tokens when applied to both relative and absolute timestamps. For both models, Chrono-BLIP and Chrono-GPT, the most evident gain of our sequence design is that interleaving frames and timestamps significantly outperform the non-interleaved settings. Our experiments demonstrate that representing time in seconds rounded to the respective nearest integer in an interleaved sequence design yields the best performance when applied to our BLIP-2 backbone. For our Chrono-GPT setting, on the other hand, relative and absolute representations are on par. For the sake of simplicity and consistency in our following experimental setup, for Chrono-GPT we use the row 6 settings.
Table 2: Comparison with state-of-the-art on QVHighlights (Lei et al., 2021a), Charades-STA (Gao et al., 2017), and ActivityNet Captions (Krishna et al., 2017). indicates QVH results are reported on the validation set.
QVHighlights Charades-STA ActivityNet
Method R1@.5 R1@.7 mAP mAP@.5 mAP@.75 R1@.5 R1@.7 R1@.5 R1@.7
Finetuned for Moment Retrieval
VLG (2021) 46.30 29.80
SeViLa (2024) 54.50 36.50
CG-DETR (2023a) 65.43 48.38 42.90 64.51 42.77 58.44 36.34
UnLoc-L (2023) 60.80 38.40 48.30 30.20
EaTR (2023) 61.36 45.79 41.74 61.86 41.91 68.47 44.92
InternVideo2-6B (2024) 71.42 56.45 49.24 70.03 48.95
BLIP-2 (frames only) 69.10 46.52 37.87 60.68 38.93 43.33 32.60 25.84 9.72
Zero-Shot
GPT-4o (frames only) 5.49 2.17 1.1 2.81 0.76 7.02 2.12 7.20 2.80
4.2 Comparison to the SOTA in Moment Retrieval
Table 3: Transfer Learning Results for moment retrieval across datasets. indicates that Momentor was trained on a subset of ActivityNet Captions.
Charades-STA ActivityNet
Method R1@.5 R1@.7 R1@.5 R1@.7
Momentor (2024) 26.60 11.60 23.00 12.40
TimeChat (2023) 32.20 13.40
VTimeLLM (2024a) 34.30 14.70
Chrono-BLIP[Uncaptioned image] 34.87 18.04 28.64 16.44
Next, we compare Chrono to the state-of-the-art (SOTA) approaches for MR following the best design choices derived from our ablations in Section 4.1. Specifically, we compare Chrono-BLIP to other finetuned baselines and Chrono-GPT to zero-shot baselines.
Firstly, in Table 2, we include a comparison to the vanilla BLIP-2 model, where we only include the frames in the context of BLIP-2 and finetune and evaluate it on each dataset, respectively. As expected, our Chrono-BLIP model significantly outperforms this baseline. Notably, for QVH, this improvement is less pronounced than on the other datasets because the videos in the dataset are always 150 seconds long, so the model can better learn to infer each frame’s temporal location since it is consistent. Nevertheless, it still performs far worse than our Chrono-BLIP model.
Compared to prior SOTA methods on QVHighlights, Chrono-BLIP significantly outperforms the previous SOTA InternVideo2 (Wang et al., 2024) on all metrics, which is a general video-language foundation model requiring expensive video pretraining. Our model improves the performance by 3.35% and 4.06% for R1@0.5 and R1@0.7, respectively.
On Charades-STA, our Chrono-BLIP model also achieves a new state-of-the-art, improving over the previous SoTA model (Wang et al., 2024) for R1@0.7. While InternVideo2 achieves good performance by extracting intermediate features and leveraging CG-DETR (Moon et al., 2023a) as a localization head, we demonstrate strong results with a fraction of the training data, smaller model size, fewer trained parameters, and while not relying on a task-specific output head. We solely finetune on the downstream task. Chrono-BLIP also outperforms the previous SoTA (Yan et al., 2023) in moment retrieval on ActivityNet Captions by 5.62% and 5.35% on R1@0.5 and R1@0.7, respectively.
Our work is most comparable to SeViLa (Yu et al., 2024), which leverages the same image-text pretrained MLLM, namely BLIP-2 (Li et al., 2023), as its backbone. Our Chrono model achieves performance that is 20.14% and 23.29% higher than SeViLa on QVHighlights in Recall@1 at thresholds of IoU=0.5 and IoU=0.7. This demonstrates the effectiveness of our novel multimodal sequence design. Our model receives contextual and temporal information and predicts the relevant moments in consecutive autoregressive forward passes. In contrast, SeViLa classifies each frame as relevant to the query or not without any contextual and temporal information about the video.
Moreover, we also evaluate our Chrono-GPT model on all three MR benchmarks in a zero-shot setting. As expected, the performance achieved by Chrono-GPT is much lower compared to methods explicitly trained for Moment Retrieval in those same datasets. However, it performs significantly better than the baseline of only providing the frames without the Chrono framework, validating that our Chrono blueprint enables moment retrieval in a zero-shot manner. This reinforces our goal of having the simplest possible approach with competitive results, and we believe can be very useful for practitioners.
Finally, to compare to more MLLM-based MR models, we evaluate Chrono-BLIP’s performance on the transfer learning setting in Table 3. Prior work (Qian et al., 2024; Ren et al., 2023; Huang et al., 2024a) perform extensive video-language instruction finetuning on multiple temporal grounding datasets and evaluate the performance on the held-out Charades-STA dataset. Although the training corpus of (Qian et al., 2024) includes a subset of ActivityNet, they also evaluate on the test respective test set, whereas our transfer learning setup involves training only on QVH and then evaluating on ActivityNet Captions. For our Charades-STA evaluation, we train on ActivityNet Captions. Although these prior works train on much larger datasets and have larger model sizes, we again achieve SoTA results.
We demonstrate the surprising effectiveness of simple but deliberate design choices for leveraging an image-text pretrained MLLM to achieve state-of-the-art results.
4.3 Grounded Video QA on NExT-GQA
Table 4: NExT-GQA test set. First four models are known to be explicitly finetuned on NExT-QA. See discussion in Section 4.3.
Method mIoU IoP@.5 A@GQA A@QA
Finetuned on Dataset
SeViLA (2024) $21.7$ 22.9 $16.6$ $68.1$
FrozenBiLM (2023) $9.6$ 23.7 $17.5$ $70.8$
Chrono-BLIP[Uncaptioned image] 28.7 24.6 19.4 73.9
Zero-Shot
Uniform 60F $21.1$ 21.1 $8.7$ $7.1$
LLoVi (2023a) $21.5$ 38.0 $26.8$
DeVi (2024) $22.3$ 37.9 $28.0$ $71.6$
Chrono-GPT[Uncaptioned image] 36.5 50.8 42.1 79.3
Next, we challenge Chrono to a new task, grounded video question answering (GQA) on the NExT-GQA (Xiao et al., 2023) benchmark, i.e., localizing a moment that is relevant to answering a question and, then even answering the question. We provide more details on the benchmark and metrics in Appendix Appendix B.
To evaluate Chrono-BLIP on GQA, we follow the localizer-answerer approach of SeViLa (Yu et al., 2024) using 60 frames for localizing the relevant moment, and then resampling 60 new frames out of that segment for answering the question. We apply our QVH-pretrained Chrono-BLIP model as the localizer and finetune a separate Flan-T5 XL LLM as our answerer. Notably, we do not finetune our localization model, thus applying it to a new domain, the domain of questions, where a relevant moment does not have to be explicitly described in the question to be relevant to answering it. We also compare to uniformly sampling 60 frames. Table 4 demonstrates the performance of Chrono-BLIP on the NExT-GQA (Xiao et al., 2023) benchmark and compares it to prior state-of-the-art models, SeViLa and FrozenBiLM (NG+) (Xiao et al., 2023).
Following the same protocol, we evaluate our Chrono-GPT model in a zero-shot setting. As ablated, our timestamp design enables strong zero-shot moment retrieval for Chrono-GPT. Paired with the strong QA abilities of a foundational model, we achieve a new state-of-the-art result on NExT-GQA over prior work (Zhang et al., 2023a; Qin et al., 2024), which also leverages GPT-4 and GPT-4o, respectively.
We detail our setup for NExT-GQA with Chrono-GPT, alongside ablations for various design decisions, in Appendix Section C.3.
4.4 Qualitative Results
Figure 3 shows qualitative results that demonstrate Chrono-BLIP’s ability to predict multiple windows (2 and 3), generalize to unseen data distributions (5), as well as some failure modes like predicting wrong segments (3) and only partially adhering to the prompt due to low image resolution (4). In (4), the query is ”A woman in long brown hair is trying on a black hat in a shop”. Chrono-BLIP predicts a moment where the woman is clearly visible in a shop but not trying on a hat. In the ground truth segment, the woman is trying on a hat but is not clearly visible since she is far away. We provide more videos with localizations in Appendix D.
@@ 5 Conclusion
We introduce Chrono[Uncaptioned image], a simple model-agnostic blueprint for providing temporal information to video-language MLLMs, thus enabling them to perform tasks such as moment retrieval and grounded video question answering. We show that originally image-text pretrained Chrono models can easily adapt to the video-language modality and greatly benefit from an input design that incorporates visual and temporal information as a simple interleaved token sequence of visual and text tokens. Specifically, finetuned Chrono-BLIP achieves state-of-the-art results on the popular moment retrieval benchmarks, whereas zero-shot Chrono-GPT achieves a new SoTA in grounded video question answering. Through our extensive ablations, we surprisingly find that our simple model-agnostic design outperforms all prior SOTA methods, which rely on aspects like task-specific model architectures, extensive video pretraining, further input signal, such as video transcripts or novel time embedding modules. We hope this versatile sequence-to-sequence design of Chrono serves as a fundamental baseline or potential best practice for encoding time in video-language models and sparks further research in leveraging image-text pretrained MLLMs for video understanding tasks.
6 Limitations and Impact
Since we leverage large-scale pretrained MLLMs, BLIP-2 (Li et al., 2023) and GPT-4o (OpenAI, 2024a), we can not rule out the presence of biases or harmful stereotypes (e.g., gender or racial biases) that these models have learned. It is conceivable that our model will detect moments that fall into the category of harmful stereotypes. We recommend applying our model with caution when adopting it in practice. Moreover, our model could be subject to misuse w.r.t some forms of surveillance.
@@ 7 Acknowledgements
The research was partially funded by a LOEWE-Start-Professur (LOEWE/4b//519/05.01.002-(0006)/94), LOEWE-Spitzen-Professur (LOEWE/ 4a//519/05.00.002-(0010)/93) and an Alexander von Humboldt Professorship in Multimodal Reliable AI, sponsored by Germany’s Federal Ministry for Education and Research.
Appendix A Appendix Overview
In the Appendix, we elaborate on further details of our work and present additional experiments. Specifically, we discuss the following:
  • Our experimental setup, detailing the benchmarks we used and our training and zero-shot setup details for reproducibility in Appendix B.
  • Ablations on the effect of the number of trainable parameters in Section C.1.
  • In Section C.2, further detailed ablations on the timestamp design for zero-shot GPT-4o on Charades-STA (Table 7) and QVHighlights (Table 8). We find that for longer videos, the effect of interleaving the timestamps greatly improves the moment localization performance, supporting our findings in the main paper.
  • The exact prompt used for querying GPT-4o (Figure 4).
  • Additional qualitative examples (Appendix D).
  • Finally, we provide a separate video montage showcasing example videos from the benchmarks and Chrono-BLIP’s predicted moment localizations and the code, respectively.
Appendix B Experimental Setup
B.1 Benchmarks
Charades-STA (Gao et al., 2017)
includes 9,848 videos with an average duration of 30.6 seconds. The dataset contains 16,128 annotations with an average moment length of 8.1 seconds and an average query length of 7.22 words. The dataset is originally divided into two splits: training (12,408) and test (3,720). To avoid overfitting on the test set during training, we designate a part of the training set as a new validation set. The videos used in the validation set are not contained in the training set. Our new dataset split consists of a split for training, validation, and testing, with 11,166, 1,242, and 3,720 annotations, respectively. For a fair comparison, after completing our ablations, we train our final model on the original training set and report numbers on the test set. We will share our split for reproducibility.
QVHighlights (Lei et al., 2021a)
is one of the most recent benchmarks and contains 10,148 videos with a duration of 150 seconds. The videos were cropped out of YouTube videos, and each video is annotated with at least one query with an average length of 11.3 words describing the relevant moment. The target windows have an average length of 24.6 seconds. The dataset is split into training, validation, and test sets with 7,218, 1,150, and 1,542 queries, respectively. This benchmark is challenging because one query can be associated with multiple moments in a video. The test set targets are withheld and thus guarantee a fair benchmark. The evaluation for the test split can only be measured through submitting the prediction to the evaluation server [^5] Eval server: https://codalab.lisn.upsaclay.fr/competitions/6937.
ActivityNet Captions (Krishna et al., 2017)
contains 20,000 videos with an average duration of 2 minutes. The dataset contains 72,000 segments, each human-annotated with a caption that includes, on average, 13.5 words. The dataset is divided into three splits, train (37,421), val_1 (17,505), and val_2 (17,031). Following (Yan et al., 2023), we use the train split for training, val_1 for validation, and val_2 for testing.
NExT-GQA (Xiao et al., 2023)
extends the NExT-QA (Xiao et al., 2021) benchmark by providing temporal grounding for the moments in the video, that are relevant for answering the question for the validation and test sets, making this a weakly-supervised dataset. The dataset contains the original training split of NExT-QA, which includes 34,132 samples. The validation set contains 3,358 questions with 3,931, meaning a sample can potentially have multiple relevant moments, but where predicting any single one of those moments is a correct prediction, i.e., one does not have to predict all relevant moments. The test set has 5,553 questions with 6,600 segments. The segments have an average duration of 7.3 and 6.7 seconds for the validation and test set, respectively.
B.1.1 Metrics
The most commonly used metrics for MR are Recall@K and mean average precision (mAP) computed under different Intersection over Union (IoU) thresholds. The Recall@K metric is defined as the percentage of the top-K predicted segments having a larger temporal IoU than the threshold with a ground truth segment. Following the recent development (Lee & Byun, 2023; Lei et al., 2021a; Moon et al., 2023a, b), we report the more challenging $R1@0.5$ and $R1@0.7$ scores, which correspond to the $Recall@1$ scores at IoU thresholds of $0.5$ and $0.7$, respectively, and the mean IoU (mIoU) score. Following these prior works, for MR on QVHighlights, we report the mAP score at IoU thresholds of $0.5$ and $0.75$ and the average mAP, respectively.
NExT-GQA (Xiao et al., 2023) introduces the metrics (i) Intersection over Prediction (IoP), which is defined as the ratio between the length of the temporal intersection between the predicted relevant moment and ground truth and the length of the predicted moment itself, and (ii) Accuracy@GQA (A@GQA) which is the accuracy of correctly answered questions that have a grounding score of $IoP>0.5$. The IoP loosens the common IoU metric by only considering the intersection over the length of the prediction instead of over the union. This metric encourages very short-moment predictions and is not concerned with a precise coverage of the entire relevant moment.
B.2 Implementation Details

### v6

@@ 4.6.2 Design of timestamps
Numerical precision of timestamps. A central question for this work is how MLLMs can best represent and reason about timestamps, as they have to be interpreted in the input and predicted in the output. We compare different options for representing the timestamps and their impact on the final downstream model performance. We explore using relative positions w.r.t. the video duration vs. representing each timestamp as absolute time. We compare representing these as decimal numbers (absolute “79.9” seconds, relative “0.40”) vs. integers (absolute “80” seconds, relative “40”). In these experiments, the format of the timestamps in model input and output is always consistent. In the case of relative positions, we post-process the output to yield an absolute value to compare to the ground truth for the final evaluation.
@@
Inspecting the Chrono-BLIP ablations in Table IV (b), we observe that representing the timestamps in decimal form results in lower performance for both absolute and relative form, see rows (1,2) vs. rows (3, 4). We hypothesize this is the case because of how decimal numbers are tokenized. Although using decimal timestamps allows for higher temporal resolution, we hypothesize that this lower performance can be attributed to limitations related to the tokenization of decimal numbers. Depending on the tokenizer, a decimal number such as “79.9” or “42.05” can be divided into a different number of tokens each, whereas integers (up to a certain number) are tokenized as a single token. Nevertheless, it is important to note that this behavior varies across tokenizers and can, therefore, be subject to variance. Further, we observe that absolute position in seconds yields better performance than the relative representation when using integer precision. This shows the importance of designing the input and outputs with the representation of the MLLM in mind.
@@
When analyzing the trends for Chrono-GPT, in Table IV (b) and, more detailed, in Table VI, we find that these trends are not as evident. We observe that there is higher variance when evaluating Chrono-GPT on the downstream tasks and that it seems, when inspecting only the first four rows, that GPT-4o prefers decimal precision. We hypothesize that GPT-4o, being a more advanced model, might be more suitable for processing numbers with decimal precision.
TABLE VI: Ablation: timestamp design for GPT-4o zero-shot on Charades-STA and QVHighlights validation sets. We compare relative vs. absolute representations, decimal vs. integer precision, and interleaving vs. appending timestamps.
Charades-STA QVHighlights
# Rep. Prec. Inter. R1@.5 R1@.7 mIoU R1@.5 R1@.7 mIoU
(1) Rel Dec 32.85 11.38 37.78 48.39 26.06 46.45
(2) Abs Dec 32.38 13.30 36.99 44.19 24.58 43.82
(3) Rel Int 34.68 15.46 38.04 30.42 14.90 33.91
(4) Abs Int 29.91 11.28 33.43 36.77 20.68 37.41
(5) Rel Int 34.86 11.20 38.44 62.84 42.19 57.91
(6) Abs Int 28.35 11.29 35.43 61.68 41.80 57.12
(7) Abs Dec 28.72 10.74 34.78 58.77 40.16 55.83
Table VI focuses on ablating the timestamp design when using GPT-4o zero-shot for moment retrieval. The QVHighlights results support the main hypothesis that interleaving timestamps between frames leads to better localization, also in the zero-shot case. The effect of interleaving is less clear with Charades-STA. We hypothesize that this difference is caused by the fact that we use only 20 frames (out of shorter videos) for Charades-STA, whereas in the QVHighlights, GPT-4o observes 60 frames (of videos up to several minutes long). This means appended timestamps are further away from the frame that they refer to, making the gap larger relative to interleaving on QVHighlights. However, in Charades-STA, the shorter distance between token frames and appended timestamps seemingly has a smaller effect. Additionally, we could hypothesize that successive frames are more in distribution for GPT-4o than a series of frames with interleaved timestamps, therefore offering some benefit to simply appending the timestamps in very short frame sequences. This illustrates how some timestamp decisions can be more beneficial if the model can be finetuned, as shown for Chrono-BLIP[Uncaptioned image]. We can also observe a difference in this zero-shot setting, which uses a stronger model. Interleaving relative integer timestamps yields better performance than using the absolute temporal representation in seconds. There are to possible explanations for this: Firstly, a relative timestamp might be more intuitive without an inherent understanding of absolute time (seconds). Secondly, the model is able to output predictions in a higher temporal resolution. Lastly, the best configurations for both datasets use integer precision rather than decimal, even if the latter offers a higher temporal resolution, which is in agreement with the results when finetuning, as shown in Table IV.
@@
Interleaving frame and time tokens. Moreover, we explore the design choice of ordering frame and time tokens in Table IV (b). In rows (1) through (4), we concatenate all frames together, followed by a concatenation of the timestamps, which are each separated by a separator token (“>>”). In rows (5) and (6), we observe the benefit of interleaving frame tokens with time tokens when applied to both relative and absolute timestamps. For both models, Chrono-BLIP and Chrono-GPT, the most evident gain of this sequence design is that interleaving frames and timestamps significantly outperforms the non-interleaved settings. The experiments demonstrate that representing time in seconds, rounded to the respective nearest integer, in an interleaved sequence design yields the best performance when applied to the BLIP-2 backbone; we use this configuration in the SOTA comparisons (Table I(a)). For the Chrono-GPT setting, on the other hand, relative and absolute representations are on par.
For simplicity and consistency in subsequent experiments, for Chrono-GPT we use the row (6) configuration: absolute integer seconds with interleaved timestamps. We adopt these best settings in the SOTA comparisons (Table I(a)).
4.6.3 Extending to different model sizes and model families
TABLE VII: Qwen2.5-VL-Instruct performance comparison with and without Chrono blueprint. We evaluate different model configurations on Charades-STA test set and measure recall at IoU thresholds of 0.5 and 0.7. For both model sizes and finetuned/zero-shot settings, the Chrono blueprint (in a grey background) consistently improves moment retrieval.
Size Finetuned Chrono R1@.5 R1@.7
3B 7.74 2.88
[…]
44.41 23.47
[…]
7B 22.07 9.73
[…]
48.20 27.98
[…]
To explore different model sizes and the ability of the Chrono blueprint to generalize to even further model architectures, we perform additional ablations using the Qwen2.5-VL [Qwen2.5-VL] family of models and report numbers on the Charades-STA validation set.
We demonstrate that the Chrono[Uncaptioned image] blueprint is useful across Qwen2.5-VL sizes and tuning regimes, as shown in Table VII. In all scenarios, finetuned/zero-shot or 3B/7B, the Chrono[Uncaptioned image] blueprint consistently improves moment retrieval.
Importantly, these experiments allow us to gauge whether M-RoPE, which encode temporal information per frame, are enough for moment retrieval. These results prove that Chrono[Uncaptioned image]’s explicit timestamp representation provides benefits beyond what M-RoPE can achieve independently, although the gap in the zero-shot case is remarkably smaller than for GPT-4o, as seen in Table I(a), which is presumably not using such positional encoding.
We find that the Qwen2.5-VL models are more sensitive to changes in the instruction/format prompt, and cannot exploit the use of chain-of-thought. This can be explained both by their smaller size, and the nature of their training, i.e. instruction following and not reasoning. Therefore, we use the simple training prompt for the Qwen2.5-VL models also in the zero-shot setting.
4.6.4 Effect of number of trainable parameters
[…]
(a) Frame count ablation.
Frames R1@.5 R1@.7
5 54.50 32.72
10 64.63 41.80
15 67.79 44.07
20 67.28 46.70
25 67.95 43.35
30 64.58 41.51
(b) LoRA rank ablation.
Rank R1@.5 R1@.7 mIoU
2 66.03 43.03 56.59
4 65.71 43.59 56.78
8 67.28 46.70 57.46
16 67.72 45.38 57.35
In Table VIII(b), we show the effect of different LoRA ranks and their resulting number of trainable parameters for Chrono-BLIP. Note that this is different from scaling the model altogether, as shown in Table VII, since we keep the total model size constant but vary the model capacity devoted to moment retrieval. For simplicity, we report results on the Charades-STA validation set with Chrono-BLIP. We observe that setting the rank to 16, i.e. training double the number of parameters does not yield any significant improvement (row 4). Moreover, training fewer parameters, i.e. 10 or 6 million, by setting the rank to 4 or 2, respectively, does degrade the performance of Chrono (rows 1 and 2).
4.6.5 Effect of number of frames
In Table VIII(a), we show the impact of the number of input frames and demonstrate the ability of the image-text pretrained MLLM to adapt to the video modality.
For Charades-STA with an average video duration of 30 seconds, using 20 frames achieves the highest R1@0.7, and offers a good efficiency-performance trade-off when considering R1@0.5 and mIoU scores.
For QVHighlights, which consists of 150-second long videos, and for ActivityNet, which on average has 120-second long videos, we have found 60 frames to perform the best, showing the ability of Chrono-BLIP to comprehend a relatively long sequence of interleaved visual and textual tokens.
@@ 5 Conclusion
We introduce Chrono[Uncaptioned image], a simple model-agnostic blueprint for representing time in multimodal large language models, thus enabling them to perform tasks such as moment retrieval and grounded video question answering. We show that originally image-text pretrained Chrono models can easily adapt to the video-language modality and greatly benefit from an input design that incorporates visual and temporal information as a simple interleaved token sequence of visual and text tokens. Specifically, finetuned Chrono-BLIP achieves state-of-the-art results on the popular moment retrieval benchmarks, whereas zero-shot Chrono-GPT achieves a new SOTA in grounded video question answering. Through extensive ablations, we find that this simple model-agnostic design outperforms all prior SOTA methods, which rely on complex modifications like task-specific model architectures, extensive video pretraining, or additional input signals, such as video transcripts or novel time embedding modules. Furthermore, we show that the Chrono[Uncaptioned image] blueprint generalizes across training setups (finetuned vs zero-shot), model architectures and sizes (Chrono-BLIP, Chrono-Qwen, Chrono-GPT) and tasks (moment retrieval, grounded video QA). We hope this versatile sequence-to-sequence design of Chrono serves as a fundamental baseline or potential best practice for encoding time in video-language models and sparks further research in leveraging image-text pretrained MLLMs for video understanding tasks.
@@ 6 Acknowledgements
The research was partially funded by a LOEWEStart-Professur (LOEWE/4b//519/05.01.002-(0006)/94), LOEWE-Spitzen-Professur (LOEWE/ 4a//519/05.00.002-(0010)/93), and an Alexander von Humboldt Professorship in Multimodal Reliable AI, sponsored by Germany’s Federal Ministry for Education and Research
Hector G. Rodriguez received the M.Sc. degree in machine learning and the B.Sc. degree in theoretical physics from University College London, London, U.K. He is currently pursuing the Ph.D. degree in computer science at the Technical University of Darmstadt, Darmstadt,Germany. His research interests include multimodal reliable AI and efficient multimodal AI.
Boris Meinardus received the B.Sc. degree in computer engineering in 2021 and the M.Sc. degree in computer science in 2024, both from the Technical University of Berlin, Berlin, Germany. He is currently a Research Scientist at Sakana AI in Tokyo, Japan. His research interests are in open-ended learning with large foundation models.
Anil Batra received the B.Tech. degree in electronics and communication from Punjab Technical University, India, in 2007, the M.Sc. degree in computer science from the International Institute of Information Technology, Hyderabad, India, in 2019 and passed the Ph.D. degree in computer science at the University of Edinburgh, Edinburgh, U.K. He is currently a Research Fellow at University of Surrey and working on Sign Language videos. His research interests include multimodal AI, video understanding, and diffusion models.
Anna Rohrbach received the M.Sc. degree in Applied Mathematics from Odesa I. I. Mechnikov National University (Ukraine) in 2010 and the Ph.D. degree in Computer Science from the Max Planck Institute for Informatics and Saarland University (Germany) in 2017. She is a Professor at the Technical University of Darmstadt (Germany), where she leads the Multimodal Grounded Learning group and is part of the Multimodal AI Lab. Previously, she was a Research Scientist at the University of California, Berkeley (USA). Her research is at the intersection of Computer Vision and Natural Language Processing, including image and video description, visual grounding, visual question answering, text-to-image synthesis, and multimodal fact-checking. Prof. Rohrbach received the German Pattern Recognition Award of the German Association for Pattern Recognition (DAGM) in 2023. She is a recipient of the prestigious €2M LOEWE Start Professorship from the State of Hesse (2024). Anna Rohrbach has served as a Reviewer and Area Chair at many top-tier machine learning conferences (CVPR, ECCV, ICCV, NeurIPS, ACL, EMNLP, ICLR) and serves as a Program Chair of ECCV 2026.
Marcus Rohrbach received the B.Sc. and M.Sc. degrees in computer science from the Technical University of Darmstadt, in 2006 and 2009, respectively, and the Ph.D. degree in computer science from the Max Planck Institute for Informatics and Saarland University, in 2014. He is a Professor at the Technical University of Darmstadt, where he leads the Multimodal Reliable AI Lab. Previously, he was a postdoctoral researcher at the University of California, Berkeley, and a Research Scientist at Facebook AI Research in Menlo Park, CA, USA. His research interests include computer vision, computational linguistics, and machine learning, with an emphasis on multimodal learning and reliable AI systems. Prof. Rohrbach holds an Alexander von Humboldt Professorship in Multimodal Reliable AI and a LOEWE Spitzen Professorship awarded by the state of Hesse, Germany.
Appendix A Implementation Details